# trajmsm: An R package for Trajectory Analysis and Causal Modeling

### v1

@@ 4.1 Features of the trajmsm package
1 $trajmsm_{i}pw$ For estimating parameters of LCGA-MSMs using IPW.
2 $trajmsm_{g}form$ For estimating parameters of LCGA-MSMs using g-computation.
3 $trajmsm_{p}ltmle$ For estimating parameters of LCGA-MSMs using pooled LTMLE.
4 $trajhrmsm_{i}pw$ For estimating parameters of LCGA-HRMSMs using IPW.
5 $trajhrmsm_{g}form$ For estimating parameters of LCGA-HRMSMs using g-computation.
6 $trajhrmsm_{p}ltmle$ For estimating parameters of LCGA-HRMSMs using pooled LTMLE.
7 $build_{t}raj$ To summarize the time-varying treatment into few groups.
@@
10 $split_{d}ata$ To split the data into multiple subsets of size $s$.
@@ 4.2.1 Identification of the trajectory groups
The next step is to identify the trajectory groups using the function $build_{t}raj$. This function allows to model the treatment and summarize the different treatment patterns into few trajectory groups. It is a wrapper of the $flexmix$ function from the R package FlexMix [12]. However, before using the $build_{t}raj$ function, the data must first be converted to long format. This can be achieved using the R base function $reshape$. Table (3) shows the data format after reshaping. It is noteworthy that the function $gendata$ can generate data directly in long format using the $format$ option.
@@
To assign individuals to a group, the $build_{t}raj$ function uses the largest posterior probability criterion (Equation 2) as suggested by Nagin [15]. The function $build_{t}raj$ currently supports binary treatments. Four arguments are required: the data in a long format with the argument Rdat, the number of trajectories with the argument $number_{t}raj$, the formula to define the polynomial form to model the trajectories and the identifier (id) of the individuals. The $build_{t}raj$ function gives as output the estimated model and the posterior probabilities of each trajectory group. For example, to fit an LCGA with $number_{t}raj$ $=3$ trajectory groups where the logit of the probability of treatment is a linear function of time:
@@
We can also extract the BIC of the estimated model from the $build_{t}raj$ object:
@@
As another example, to fit a model with $number_{t}raj$ $=3$ trajectory groups of quadratic order, the following code can be used:
@@ 4.2.3 Estimating the parameters of an LCGA-MSM using IPW
The function $trajmsm_{i}pw$ estimates the parameters of an LCGA-MSM using IPW. It provides the estimates, standard errors, confidence intervals, and p-values for the weighted regression of the outcome of interest on the trajectory groups. The function $trajmsm_{i}pw$ accounts for potential time-varying confounders as well as treatment-confounder feedback when estimating the treatment effect.
@@ 4.2.4 Estimating the parameters of an LCGA-MSM with g-computation
As discussed in Section 2.3, estimation with g-computation in the context of LCGA-MSM involves four key steps: (1) estimating the counterfactual means, (2) assigning treatment patterns to the trajectory groups, (3) regressing the counterfactual means onto the trajectory groups, and (4) performing inference using a bootstrapping approach to quantify uncertainty. The function $trajmsm_{g}form$ facilitates this process by taking the following inputs:
@@ 4.2.5 Estimating the parameters of an LCGA-MSM using pooled LTMLE
Estimation of the parameters of an LCGA-MSM with pooled LTMLE is achieved using the function $trajmsm_{p}ltmle$. This function takes the same arguments as the function $trajmsm_{g}form$, except that we do not need to specify the number of repetitions since standard errors are not computed using bootstrap but rather the influence functions. Results are presented in Table (8).
@@ 4.3 LCGA in the context of HRMSM
To begin, we generate a dataset that includes a sample of individuals over a specified follow-up period with baseline characteristics (e.g., age, sex), treatment variables (e.g., statins), and covariates (e.g., hypertension, BMI). Afterward, we segment the data into multiple time intervals using the $split_{d}ata$ function.The function $split\_data$ converts the data into the appropriate format for LCGA-HRMSM analysis. The inputs for this function are:
@@ 4.4 Estimation of the parameters of an LCGA-HRMSM
  • $class_{v}ar$: Name of the trajectory group variable.
The reference group is chosen based on the lowest observed treatment proportion and is directly handled by the algorithm. The output includes the results of the trajectory group estimation, the proportion of the treatment across all groups and time intervals, and the estimated parameters of the LCGA-HRMSM along with their standard errors and 95% confidence intervals. In the case of g-computation, the vector of coefficients resulting from the block-bootstrapping is also provided. Currently, the package supports a binary time-dependent outcome and a binary time-varying exposure. For all the functions, the names of the variables should be specified. The $var_{c}ov$ argument indicates the time-varying variables.
@@ 4.4.1 Estimating the parameters of an LCGA-HRMSM using IPW
The function $trajhrmsm_{i}pw$ estimates the parameters of LCGA-HRMSMs using IPW. Below is an example where the $total_{f}ollowup$ $=8$ time points are split in three subsets of size $ntimes_{i}nterval$ $=6$ each (1-6, 2-7, 3-8) and $number_{t}raj$ $=3$ groups are formed.
@@
In Table (10), the observed mean of the treatment per trajectory group across all periods is shown ; such a result can be displayed using $res_{t}rajhrmsm_{i}pw[[3]]$.
@@
In Table (11), results of the estimation are displayed using the line of code: $reshrmsm_{i}pw[[2]]$.
@@ 6.0.1 Estimating parameters of LCGA-HRMSM using g-computation
To estimate the parameters of LCGA-HRMSMs using g-computation, the function $trajhrmsm_{g}form$ is used. Results of the estimation are presented in Table (12).
@@ 6.0.2 Estimating parameters of LCGA-HRMSM using pooled LTMLE
Results of the estimation are presented in Table (13). Results when using the function $trajhrmsm_{p}ltmle$:

### v2

@@ 4.1 Features of the trajmsm package
1 $trajmsm\_ipw$ For estimating parameters of LCGA-MSMs using IPW.
2 $trajmsm\_gform$ For estimating parameters of LCGA-MSMs using g-computation.
3 $trajmsm\_pltmle$ For estimating parameters of LCGA-MSMs using pooled LTMLE.
4 $trajhrmsm\_ipw$ For estimating parameters of LCGA-HRMSMs using IPW.
5 $trajhrmsm\_gform$ For estimating parameters of LCGA-HRMSMs using g-computation.
6 $trajhrmsm\_pltmle$ For estimating parameters of LCGA-HRMSMs using pooled LTMLE.
7 $build\_traj$ To summarize the time-varying treatment into few groups.
@@
10 $split\_data$ To split the data into multiple subsets of size $s$.
@@ 4.2.1 Identification of the trajectory groups
The next step is to identify the trajectory groups using the function $build\_traj$. This function allows to model the treatment and summarize the different treatment patterns into few trajectory groups. It is a wrapper of the $flexmix$ function from the R package FlexMix [12]. However, before using the $build\_traj$ function, the data must first be converted to long format. This can be achieved using the R base function $reshape$. Table (3) shows the data format after reshaping. It is noteworthy that the function $gendata$ can generate data directly in long format using the $format$ option.
@@
To assign individuals to a group, the $build\_traj$ function uses the largest posterior probability criterion (Equation 2) as suggested by Nagin [15]. The function $build\_traj$ currently supports binary treatments. Four arguments are required: the data in a long format with the argument Rdat, the number of trajectories with the argument $number\_traj$, the formula to define the polynomial form to model the trajectories and the identifier (id) of the individuals. The $build\_traj$ function gives as output the estimated model and the posterior probabilities of each trajectory group. For example, to fit an LCGA with $number\_traj$ $=3$ trajectory groups where the logit of the probability of treatment is a linear function of time:
@@
We can also extract the BIC of the estimated model from the $build\_traj$ object:
@@
As another example, to fit a model with $number\_traj$ $=3$ trajectory groups of quadratic order, the following code can be used:
@@ 4.2.3 Estimating the parameters of an LCGA-MSM using IPW
The function $trajmsm\_ipw$ estimates the parameters of an LCGA-MSM using IPW. It provides the estimates, standard errors, confidence intervals, and p-values for the weighted regression of the outcome of interest on the trajectory groups. The function $trajmsm\_ipw$ accounts for potential time-varying confounders as well as treatment-confounder feedback when estimating the treatment effect.
@@ 4.2.4 Estimating the parameters of an LCGA-MSM with g-computation
As discussed in Section 2.3, estimation with g-computation in the context of LCGA-MSM involves four key steps: (1) estimating the counterfactual means, (2) assigning treatment patterns to the trajectory groups, (3) regressing the counterfactual means onto the trajectory groups, and (4) performing inference using a bootstrapping approach to quantify uncertainty. The function $trajmsm\_gform$ facilitates this process by taking the following inputs:
@@ 4.2.5 Estimating the parameters of an LCGA-MSM using pooled LTMLE
Estimation of the parameters of an LCGA-MSM with pooled LTMLE is achieved using the function $trajmsm\_pltmle$. This function takes the same arguments as the function $trajmsm\_gform$, except that we do not need to specify the number of repetitions since standard errors are not computed using bootstrap but rather the influence functions. Results are presented in Table (8).
@@ 4.3 LCGA in the context of HRMSM
To begin, we generate a dataset that includes a sample of individuals over a specified follow-up period with baseline characteristics (e.g., age, sex), treatment variables (e.g., statins), and covariates (e.g., hypertension, BMI). Afterward, we segment the data into multiple time intervals using the $split\_data$ function.The function $split\_data$ converts the data into the appropriate format for LCGA-HRMSM analysis. The inputs for this function are:
@@ 4.4 Estimation of the parameters of an LCGA-HRMSM
  • $class\_var$: Name of the trajectory group variable.
The reference group is chosen based on the lowest observed treatment proportion and is directly handled by the algorithm. The output includes the results of the trajectory group estimation, the proportion of the treatment across all groups and time intervals, and the estimated parameters of the LCGA-HRMSM along with their standard errors and 95% confidence intervals. In the case of g-computation, the vector of coefficients resulting from the block-bootstrapping is also provided. Currently, the package supports a binary time-dependent outcome and a binary time-varying exposure. For all the functions, the names of the variables should be specified. The $var\_cov$ argument indicates the time-varying variables.
@@ 4.4.1 Estimating the parameters of an LCGA-HRMSM using IPW
The function $trajhrmsm\_ipw$ estimates the parameters of LCGA-HRMSMs using IPW. Below is an example where the $total\_followup$ $=8$ time points are split in three subsets of size $ntimes\_interval$ $=6$ each (1-6, 2-7, 3-8) and $number\_traj$ $=3$ groups are formed.
@@
In Table (10), the observed mean of the treatment per trajectory group across all periods is shown ; such a result can be displayed using $res\_trajhrmsm\_ipw[[3]]$.
@@
In Table (11), results of the estimation are displayed using the line of code: $reshrmsm\_ipw[[2]]$.
@@ 6.0.1 Estimating parameters of LCGA-HRMSM using g-computation
To estimate the parameters of LCGA-HRMSMs using g-computation, the function $trajhrmsm\_gform$ is used. Results of the estimation are presented in Table (12).
@@ 6.0.2 Estimating parameters of LCGA-HRMSM using pooled LTMLE
Results of the estimation are presented in Table (13). Results when using the function $trajhrmsm\_pltmle$: